OpenStudy (clinton):

at what altitude above the earth's surface would the acceleration due to gravity be 4.9ms-2 assuming the mean radius of the earth is 6.4 EXP 6 meters and acceleration due to gravity 9.8ms-2 on the surface of the earth

OpenStudy (mos1635):

go=G*M/R^2 on serface g=G*M/r^2 up high divide them, find r h=r-R

OpenStudy (mos1635):